Brigham and Houston just paid a dividend of $2.00 a share. The dividend is expected to grow at 25% a year for the next 3 years and then at a rate of 6% a year thereafter.

What is the expected dividend per share for year 3?

Group of answer choices
Answer #1

Year 1 dividend = 2 + (1 + 25%) = 2.5

Year 2 dividend = 2.5 + (1 + 25%) = 3.125

Year 3 dividend = 3.125 + (1 + 25%) = 3.9063

Expected dividend in year 3 = 3.9063 or 3.91